# IPPP/25/40

### v1

@@ I Introduction
The power spectrum of primordial density fluctuations, $\mathcal{P}(k)$, is a key quantity in cosmology. It describes the initial conditions of the Universe right after the big bang, tiny perturbations in an almost homogeneous plasma that eventually evolved into all large-scale cosmological structures that we can see today. Exquisite observations of the cosmic microwave background radiation (CMB) [1] and large-scale structure, in particular through the Lyman-$\alpha$ forest [2, 3], tightly constrain $\mathcal{P}(k)$ on length scales larger than $\sim 1$ Mpc. These observations are consistent with the simplest models of cosmological inflation [4], which represents a remarkable success for our leading explanation of how those primordial perturbations were seeded initially. $\mathcal{P}(k)$ is thus not only crucial for understanding the universe at large scales, but also contains invaluable information about fundamental new physics, connected to much higher energy scales than what is accessible by earth-bound experiments.
Fully exploiting this potential would require to extend our knowledge of $\mathcal{P}(k)$ to sub-Mpc scales, but the heap of information that may be buried there remains largely inaccessible at present. At the smallest scales, in particular, the non-observation of primordial black holes (PBHs) provides our only probe of $\mathcal{P}(k)$; this leads to rather mild constraints, about seven orders of magnitude weaker than what is observed at large scales [5, 6]. Ultracompact minihalos (UCMHs) form much more frequently from rare density fluctuations in the early universe than PBHs [7, 8, 9]. Assuming that dark matter (DM) can pair-annihilate into standard model particles, like in the case of the much-discussed weakly interacting massive particle (WIMP) explanation of DM [10, 11], this translates into very stringent small-scale constraints on $\mathcal{P}(k)$ [12, 9, 13, 14, 15, 16, 17, 18]. Future pulsar timing array observations of UCMH transients might constrain $\mathcal{P}(k)$ independent of this assumption [19, 20]. Later forming DM halos may also lead to various constraints at sub-Mpc scales, subject to assumptions about the late-time evolution and properties of these objects [21, 22, 23, 24, 25, 26, 27]. Finally, spectral distortions of the CMB [28] and the induced gravitational wave background [29, 30] provide complementary constraints directly on the linear, or only mildly non-linear, power spectrum at scales down to roughly one pc.
@@
Here we combine the idea of using UCMHs to constrain the primordial power spectrum with these recent developments in the more general EDO context. We present new and significantly extended limits on EDOs, and apply them to the more specific case of UCMHs. In combination, this allows us to improve existing constraints on $\mathcal{P}(k)$ by up to an order of magnitude, over a wide range of sub-Mpc scales. Such tightened small-scale constraints will not only be decisive in further narrowing down the microphysics governing inflation (see, e.g., Ref. [37]), but also put highly relevant bounds on other mechanisms generating excess power on small scales, like dark sector phase transitions [18].
@@ II UCMH formation and properties
where $a$ is the scale factor, $M_{\rm Pl}$ the Planck mass, $\rho_{\chi}$ the DM density and we introduced $\Omega_{\chi}\equiv\rho_{\chi}/\rho_{c}=0.264$ [1], with $\rho_{c}$ being the critical density. It was originally argued that accretion would enhance $M_{\rm UCMH}$ by another factor of $a_{\rm eq}/a$ after equality, but such a growth could not consistently be reproduced in simulations [38] and we hence conservatively neglect it for most of our work.
@@ III Updated EDO limits
Recently, gravitational constraints on generic EDOs have been derived from the non-observation of micro-lensing events [33, 31], gravitational wave production [34], the evaporation of wide binaries [49, 50], dynamical heating of dwarf galaxies [51], and heating of the CMB due to baryonic accretion [35]. A repository of current constraints on EDOs with different mass profiles is provided in Ref. [36]. In this work, we extend previous bounds on EDOs from accretion and introduce a new constraint from microlensing of the MACS J1149 LS1 “Icarus” star.
@@
Observations of highly magnified background stars behind galaxy clusters yield a novel independent bound on EDOs. In favourable lensing configurations, where a distant star is strongly amplified by a cluster caustic, the magnification hinges on the smoothness of the cluster’s mass profile. If DM consists of compact objects, it fragments the caustic into a network of microcaustics, modifying the lensing signal. The lensed observation of the Icarus star has so far only been used to constrain the DM fraction in PBHs [53, 54], which we explain in the Appendix. In an extended macrolens, the effective Einstein radius of a point lens at distance $D_{L}$ and Einstein angle $\theta_{E}$ is $\bar{R}_{E}=\sqrt{\mu_{t}}\,\theta_{E}\,D_{L},$ which for MACS J1149 LS1 is boosted by at least $\mu_{t}\sim 15.4$ [53]. Lenses behave as point-like when their physical size satisfies [31, 33]
$R_{\rm 90}\lesssim\bar{R}_{E}=\sqrt{\mu_{t}}\,\theta_{E}\,D_{L}\approx 2.0\times 10^{6}\bigl{(}M/M_{\odot}\bigr{)}^{1/2}R_{\odot}\,,$ (2)
@@
Figure 1: Present bounds on the DM fraction $f_{\rm DM}$ comprised of EDOs (blue), with fixed radii as stated in the legend, and UCMHs (red). Constraints deriving from lensing observations of the Icarus star and from CMB heating due to baryonic accretion are described in detail in the text. We further apply Galactic microlensing [31, 33] and wide binary evaporation constraints [49].
@@ IV UCMH constraints on the power spectrum
which depends only very weakly on the upper integration boundary $\delta^{\rm max}_{\chi}\simeq\mathcal{O}(1)\gg\delta^{\rm min}_{\chi}$ corresponding to the formation of a PBH rather than an UCMH. Here, the mass variance $\sigma_{\chi,H}$ at horizon crossing of the scale $R$ is computed by convolving the power spectrum with a top-hat window function [56, 9]. For scales $k_{R}=1/R$ entering during radiation domination, it takes the form
@@
Neglecting the accretion of baryonic matter, the final UCMH abundance is given as $\Omega_{\rm UCMH}=\Omega_{\rm DM}\beta$. Comparing this to the limits above leads to the constraints on $\mathcal{P}_{\mathcal{R}}(k)$ shown in Fig. 2. It is clear that CMB accretion, wide binary evaporation and the Icarus star are highly complementary, leading to competitive bounds for scales up to a factor of $\sim 10^{4}$ smaller than what can be directly measured by large-scale structure observations.
Further constraints.—
For this range of scales, it has previously been pointed out that also the lack of Compton-$y$ and $\mu$ distortions of the CMB black-body spectrum constrains scalar perturbations. We follow Ref. [57] in computing both $y$ and $\mu$ from a given power spectrum, assuming again $\mathcal{P}_{\mathcal{R}}(k)\simeq const.$ around the pivot scale $k_{R}$. For concreteness, we assume that $\mathcal{P}_{\mathcal{R}}(k)$ is nonzero only for the range $1/3\lesssim k/k_{R}\lesssim 3$. Comparing the result to the $2\sigma$ limits of $y<1.5\times 10^{-5}$ and $\mu<9\times 10^{-5}$ based on FIRAS observations [58] then leads to the constraints shown as red lines in Fig. 2. We caution, however, that $\mu$ and $y$ in general receive contributions from a significantly larger range of scales than $\sigma_{\chi,H}$ in Eq. (4), which necessarily makes any comparison with UCMH limits somewhat model-dependent (see Appendix for a more detailed discussion).
Another relevant constraint that is independent of UCMH formation derives from the fact that scalar perturbations produce a stochastic gravitational wave background (SGWB) at second order in perturbation theory [59, 60]. The computation of the resulting power spectrum in GWs, $\mathcal{P}_{h}$, involves highly oscillatory integrals. We perform those semi-analytically by adopting an averaging procedure as advocated in Refs. [61, 62], noting that the innermost integral is dominated by contributions from modes well inside the horizon, and by assuming $\mathcal{P}_{\mathcal{R}}(k)$ to be scale-invariant over the same range of scales as for CMB distortions. We then compute the resulting fractional energy density of gravitational waves today, $\Omega_{\rm GW}h^{2}(k)=(\Omega_{r}h^{2})\frac{1}{24}({k}/aH)^{2}\mathcal{P}_{h}(k,t)$, where the r.h.s. is evaluated at any time $t_{k}\lesssim t\lesssim t_{\rm eq}$ and the factor of $\Omega_{r}h^{2}=4.15\cdot 10^{-5}$ [58] accounts for the red-shifting of $\Omega_{\rm GW}$ during matter domination. We display the resulting limits on $\mathcal{P}_{\mathcal{R}}(k)$ as black lines in Fig. 2, resulting from the requirement of $\Omega_{\rm GW}h^{2}$ not exceeding the NANOGrav data [63] at 99% C.L in either of the frequency bins.
@@ V Discussion
We have demonstrated that the formation of UCMHs leads to highly competitive constraints on $\mathcal{P}_{\mathcal{R}}(k)$ down to scales $k\sim 10^{5}$ kpc. Remarkably, these constraints do not depend on interactions between DM and ordinary matter and hence apply to any collision-less and sufficiently cold DM candidate (in particular, these constraints only apply below the scale of free streaming or kinetic decoupling [65], $k<k_{D}$, for DM particles that have been in thermal equilibrium in the early universe). We stress that we have followed a very conservative approach by restricting our analysis to objects that have fully collapsed very early, before a redshift of $z_{c}=1200$. Notably, this considerably simplifies the analysis as it largely avoids complications due to potential changes in the inner density profile or tidal disruption affecting the survival probability at later formation times.
@@
Still, even significantly smaller initial density contrasts than considered here would imply perturbations that enter the non-linear regime well before the onset of standard structure formation. While UCMHs forming at $z_{c}\lesssim 1000$ no longer efficiently contribute to the CMB accretion bounds, both wide binary evaporation and the Icarus observation therefore generally lead to significantly stronger limits than what we reported in Fig. 2. Due to the aforementioned effects, however, such more aggressive limits should be considered less robust. To a much lesser degree, our constraints are affected by other assumptions about the UCMH properties. Taking into account a potential mass growth due to baryonic accretion, in particular, would lead to somewhat stronger bounds, while assuming a shallower profile (as expected for later collapse) would result in slightly weaker constraints. For a more detailed discussion of theses aspects, we refer to the Appendix.
@@
Before concluding, let us briefly discuss further, complementary probes of the small-scale power spectrum. To start with, microlensing is effective when the physical size of the lens is smaller than the Einstein radius $R_{E}\leq R_{E,\text{max}}=\sqrt{{GMD_{S}}/{c^{2}}}$, where $D_{S}$ is the distance to the source and $M$ the lens mass, here assumed to be an UCMH. For Galactic microlensing, this implies $M>10^{{17}}\rm M_{\odot}$, which is far above the masses to which surveys are sensitive. This explains why there are no UCMH microlensing constraints in Figs. 1 and 2 but, if a greater UCMH compactness is assumed, future microlensing observations can lead to constraints on $\mathcal{P}(k)$ [66]. In contrast, the larger source distance and effective boost in $R_{E}$ make highly magnified stars such as Icarus an effective probe. Moreover, future observations of gravitational wave lensing with both sources and lenses at cosmological distances may also lead to constraints on UCMH [67, 68]. Astrometric microlensing has also been suggested as a future probe [69], as this can be detectable even if the magnification is small.
@@
VI Conclusions
While PBHs currently remain the only probe of the smallest cosmological scales, with $k\gtrsim 10^{-6}$ Mpc-1, there are several independent observations that put much more stringent constraints on the primordial power spectrum at slightly larger scales. Here we pointed out that very early forming UCMHs can be used to significantly tighten these bounds, in a way that is both robust and largely independent of the nature of DM. Notably, this also provides complementary bounds on PBH DM at these scales. Future optical and gravitational wave lensing observations have the potential of even further improvement. Such probes of the small-scale primordial power spectrum constitute a crucial step towards understanding the microphysics governing comic inflation, as well as more exotic processes that may generate excess power at these scales.
@@ VI Conclusions
TB gratefully acknowledges support through a FRIPRO grant of the Norwegian Research Council (project ID 353561 ‘DarkTurns’). DC and SSM are supported by the STFC under Grant No. ST/T001011/1. TB and SSM are grateful to the Mainz Institute for Theoretical Physics (MITP) of the Cluster of Excellence PRISMA+ (Project ID 390831469), for its hospitality and partial support during the DMLAND2024 workshop where the idea for this work was born.
Appendix A UCMH formation

### v2

@@ I Introduction
The power spectrum of primordial density fluctuations, $\mathcal{P}(k)$, is a key quantity in cosmology. It describes the initial conditions of the Universe right after the big bang, tiny perturbations in an almost homogeneous plasma that eventually evolved into all large-scale cosmological structures that we can see today. Exquisite observations of the cosmic microwave background radiation (CMB) [1] and large-scale structure, in particular through the Lyman-$\alpha$ forest [2, 3], tightly constrain $\mathcal{P}(k)$ on length scales larger than $\sim 1$ Mpc. These observations are consistent with the simplest models of cosmological inflation [4, 5], which represents a remarkable success for our leading explanation of how those primordial perturbations were seeded initially. $\mathcal{P}(k)$ is thus not only crucial for understanding the universe at large scales. It also contains invaluable information about fundamental new physics, connected to much higher energy scales than what is accessible by earth-bound experiments.
Fully exploiting this potential would require to extend our knowledge of $\mathcal{P}(k)$ to sub-Mpc scales, but the heap of information that may be buried there remains largely inaccessible at present. At the smallest scales, in particular, the non-observation of primordial black holes (PBHs) provides our only probe of $\mathcal{P}(k)$; this leads to rather mild constraints, about seven orders of magnitude weaker than what is observed at large scales [6, 7, 8]. Ultracompact minihalos (UCMHs) form much more frequently from rare density fluctuations in the early universe than PBHs [9, 10, 11]. Assuming that dark matter (DM) can pair-annihilate into standard model particles, like in the case of the much-discussed weakly interacting massive particle explanation of DM [12, 13], this translates into very stringent small-scale constraints on $\mathcal{P}(k)$ [14, 15, 11, 16, 17, 18, 19, 20, 21]. Future pulsar timing array (PTA) observations of UCMH transients might constrain $\mathcal{P}(k)$ independent of this assumption [22, 23]. Later forming DM halos may also lead to various constraints at sub-Mpc scales, subject to assumptions about late-time evolution and properties of these objects [24, 25, 26, 27, 28, 29, 30, 31]. Finally, spectral distortions of the CMB [32] and the induced gravitational wave background [33, 34] provide complementary constraints directly on the linear, or only mildly non-linear, power spectrum at scales down to roughly one pc.
@@
Here we combine the idea of using UCMHs to constrain the primordial power spectrum with these recent developments in the more general EDO context. We present new and significantly extended limits on EDOs, and apply them to the more specific case of UCMHs. In combination, this allows us to improve existing constraints on $\mathcal{P}(k)$ by up to an order of magnitude, over a wide range of sub-Mpc scales. Such tightened small-scale constraints will not only be decisive in further narrowing down the microphysics governing inflation [41, 42, 43, 44, 45], but also put highly relevant bounds on other mechanisms generating excess power on small scales, like dark sector phase transitions [21].
@@ II UCMH formation and properties
where $a$ is the scale factor, $M_{\rm Pl}$ the Planck mass, $\rho_{\chi}$ the DM density and we introduced $\Omega_{\chi}\equiv\rho_{\chi}/\rho_{c}=0.264$ [1], with $\rho_{c}$ being the critical density. It was originally argued that accretion would enhance $M_{\rm UCMH}$ by another factor of $a_{\rm eq}/a$ after equality, but such a growth could not consistently be reproduced in simulations [46] and we hence conservatively neglect it for most of our work (see Appendix for a more detailed discussion).
@@ III Updated EDO limits
Recently, gravitational constraints on generic EDOs have been derived from the non-observation of micro-lensing events [37, 35], gravitational wave production [38], the evaporation of wide binaries [58, 59], dynamical heating of dwarf galaxies [60], and heating of the CMB due to baryonic accretion [39]. A repository of current constraints on EDOs with different mass profiles is provided in Ref. [40]. Here we extend bounds from accretion and introduce a new constraint from microlensing of the MACS J1149 LS1 “Icarus” star.
@@
Observations of highly magnified background stars behind galaxy clusters yield a novel independent bound on EDOs (see Appendix) In favourable lensing configurations, where a distant star is strongly amplified by a cluster caustic, the magnification hinges on the smoothness of the cluster’s mass profile. If DM consists of compact objects, it fragments the caustic into a network of microcaustics, modifying the lensing signal. Lensed observations of MACS J1149 LS1 have so far only been used to constrain the DM fraction in PBHs [62, 63]. In an extended macrolens, the effective Einstein radius of a point lens at distance $D_{L}$ and Einstein angle $\theta_{E}$ is $\bar{R}_{E}=\sqrt{\mu_{t}}\,\theta_{E}\,D_{L},$ which for MACS J1149 LS1 is boosted by at least $\mu_{t}\sim 15.4$ [62]. Lenses behave as point-like for a physical size [35, 37]
$R_{\rm 90}\lesssim\bar{R}_{E}\approx 2.0\times 10^{6}\bigl{(}M/M_{\odot}\bigr{)}^{1/2}R_{\odot}\,$, (2)
@@
Figure 1: Present bounds on the DM fraction $f_{\rm DM}$ comprised of EDOs (blue), with fixed radii as stated in the legend, and UCMHs (red). Constraints deriving from lensing observations of the Icarus star (MACS J1149 LS1) and from CMB heating due to baryonic accretion are described in detail in the text. We further apply Galactic microlensing [35, 37] and wide binary evaporation constraints [58].
@@ IV UCMH constraints on the power spectrum
which depends only very weakly on the upper integration limit $\delta^{\rm max}_{\chi}\simeq\mathcal{O}(1)\gg\delta^{\rm min}_{\chi}$ corresponding to the formation of a PBH rather than an UCMH. Here, the mass variance $\sigma_{\chi,H}$ at horizon crossing of the scale $R$ is computed by convolving the power spectrum with a top-hat window function [65, 11]. For scales $k_{R}=1/R$ entering during radiation domination, it takes the form
@@
Neglecting the accretion of baryonic matter, the final UCMH abundance is given as $\Omega_{\rm UCMH}=\Omega_{\chi}\beta$. Requiring this to be smaller than the limits on $f_{\rm DM}$ reported above leads to the constraints on $\mathcal{P}_{\mathcal{R}}(k)$ shown in Fig. 2. Note that for our default choice of $z_{c}=1200$, the Icarus constraint does not apply for $k\gtrsim 3\cdot 10^{4}\,{\rm Mpc}^{-1}$, cf. Fig. 1, because $R^{(90)}_{\rm UCMH}$ then becomes smaller than the Einstein radius of the UCMH. For these scales, we instead adopt the smallest $z_{c}>1200$ that still results in a constraint at a given $k$ (see also Fig. 3 in Appendix).
V Further constraints
For this range of scales, it has previously been pointed out that also the lack of Compton-$y$ and $\mu$ distortions of the CMB black-body spectrum constrains scalar perturbations. We follow Ref. [66] in computing $y$ and $\mu$, assuming again $\mathcal{P}_{\mathcal{R}}(k)\simeq const.$ around the pivot scale $k_{R}$. For concreteness, we assume that $\mathcal{P}_{\mathcal{R}}(k)$ is nonzero only for the range $1/3\lesssim k/k_{R}\lesssim 3$. Comparing the result to the $2\sigma$ limits of $y<1.5\times 10^{-5}$ and $\mu<9\times 10^{-5}$ from FIRAS observations [67] then leads to the constraints shown as red lines in Fig. 2. We caution, however, that $\mu$ and $y$ in general receive contributions from a significantly larger range of scales than $\sigma_{\chi,H}$ in Eq. (4), which necessarily makes any comparison with UCMH limits somewhat model-dependent (see Appendix for a more detailed discussion).
Another relevant constraint that is independent of UCMH formation derives from the fact that scalar perturbations produce a stochastic gravitational wave background at second order in perturbation theory [68, 69]. We compute the resulting power spectrum in gravitational waves, $\mathcal{P}_{h}$, semi-analytically by adopting an averaging procedure as advocated in Refs. [70, 71], noting that the innermost integral is dominated by contributions from modes well inside the horizon. We assume $\mathcal{P}_{\mathcal{R}}(k)$ to be scale-invariant over the same range of scales as for CMB distortions. We then compute the resulting fractional energy density of gravitational waves today, $\Omega_{\rm GW}h^{2}(k)=(\Omega_{r}h^{2})\frac{1}{24}({k}/aH)^{2}\mathcal{P}_{h}(k,t)$, where the r.h.s. is evaluated at any time $t_{k}\lesssim t\lesssim t_{\rm eq}$ and the factor of $\Omega_{r}h^{2}=4.15\cdot 10^{-5}$ [67] accounts for the red-shifting of $\Omega_{\rm GW}$ during matter domination. We display the resulting limits on $\mathcal{P}_{\mathcal{R}}(k)$ as black lines in Fig. 2, resulting from the requirement of $\Omega_{\rm GW}h^{2}$ not exceeding the NANOGrav data [72] at 99% C.L in either of the frequency bins.
@@ VI Discussion
We have demonstrated that the formation of UCMHs leads to highly competitive constraints on $\mathcal{P}_{\mathcal{R}}(k)$ down to scales $k\sim 10^{5}$ kpc. Remarkably, these constraints do not depend on interactions between DM and ordinary matter and hence apply to any collision-less and sufficiently cold DM candidate (in particular, these constraints only apply below the scale of free streaming or kinetic decoupling [74], $k<k_{D}$, for DM particles that have been in thermal equilibrium in the early universe – though it should be stressed that the complementary constraints shown in Fig. 2 also assume that there are no additional physical processes washing out $\mathcal{P}_{\mathcal{R}}(k)$ on sub-horizon scales ). We emphasize that we have followed a very conservative approach by restricting our analysis to objects that have fully collapsed very early, before a redshift of $z_{c}=1200$. Notably, this considerably simplifies the analysis as it largely avoids complications due to potential changes in the inner density profile or tidal disruption affecting the survival probability at later formation times.
@@
Still, even significantly smaller initial density contrasts than considered here would imply perturbations that enter the non-linear regime well before the onset of standard structure formation. While UCMHs forming at $z_{c}\lesssim 1000$ no longer efficiently contribute to the CMB accretion bounds, both wide binary evaporation and the Icarus observation therefore generally lead to significantly stronger limits than what we reported in Fig. 2. Due to the aforementioned effects, however, such more aggressive limits should be considered less robust. To a much lesser degree, our constraints are affected by other assumptions about the UCMH properties. Taking into account a potential mass growth due to baryonic accretion, in particular, would lead to somewhat stronger bounds, while assuming a shallower profile (as expected for later collapse) would result in slightly weaker constraints. For a more detailed discussion of these aspects, we refer to the Appendix.
@@
Before concluding, let us briefly discuss further, complementary probes of the small-scale power spectrum. To start with, microlensing is effective when the physical size of the lens is smaller than the Einstein radius $R^{(90)}_{\rm UCMH}\leq R_{E,\text{max}}=\sqrt{{GMD_{S}}/{c^{2}}}$, where $D_{S}$ is the distance to the source and $M$ the lens mass. For Galactic microlensing, this formally implies $M>10^{{17}}\,\rm M_{\odot}$, i.e. far above the entire Galaxy’s mass. This explains why there are no UCMH microlensing constraints in Figs. 1 and 2 but, if a greater UCMH compactness is assumed, future microlensing observations can lead to constraints on $\mathcal{P}(k)$ [75]. In contrast, the larger source distance and effective boost in $R_{E}$ make highly magnified stars such as Icarus an effective probe. Future observations of gravitational wave lensing with both sources and lenses at cosmological distances may also lead to constraints on UCMH [76, 77]. Astrometric microlensing, finally, can be detectable even if the magnification is small and may provide a further future probe [78].
@@
VII Conclusions
While PBHs currently remain the only probe of the smallest cosmological scales, with $k\gtrsim 10^{5}$ Mpc-1, there are several independent observations that put much more stringent constraints on the primordial power spectrum at slightly larger scales. Here we pointed out that very early forming UCMHs can be used to significantly tighten these bounds, in a way that is both robust and largely independent of the nature of DM. Notably, this also provides complementary bounds on PBH DM at these scales, essentially excluding the possibility of even sub-dominant populations, with $f_{\rm PBH}\ll 1$, that formed from the collapse of initial density perturbations at scales $k\lesssim 10^{5}\,{\rm Mpc}^{-1}$ (corresponding to $M_{\rm PBH}\gtrsim 10^{2}\,M_{\odot}$). Future optical and gravitational wave lensing observations have the potential of even further improvement. Such probes of the small-scale primordial power spectrum constitute a crucial step towards understanding the microphysics governing cosmic inflation, as well as more exotic processes that may generate excess power at these scales.
@@ VII Conclusions
We thank Kai Schmidt-Hoberg for very useful discussions, leading to improvements of an initial version of the draft. TB gratefully acknowledges support through a FRIPRO grant of the Norwegian Research Council (project ID 353561 ‘DarkTurns’). DC and SSM are supported by the STFC under Grant No. ST/T001011/1. TB and SSM are grateful to the Mainz Institute for Theoretical Physics (MITP) of the Cluster of Excellence PRISMA+ (Project ID 390831469), for its hospitality and partial support during the DMLAND2024 workshop where the idea for this work was born.
Appendix A Appendix A: UCMH formation